I can not figure this out...need help
Linear Equations:Clearing Fractions and Decimals
1/6X+7/10=X

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

1/6X+7/10=X
1/6X-X=-7/10
-5/6X=-7/10
5/6X=7/10 CROSS MULTIPLY WE GET
50X=42
X=42/50
X=21/25

Answer by stanbon(75887)

(1/6)X + (7/10) = X
Subtract (1/6)x from both sides to get:
(5/6)x = 7/10
x= 42/50
x= 21/25
-------
Cheers,
Stan H.